### v1

# ToolACE-R: Tool Learning with Adaptive Self-Refinement

Abstract
Tool learning, which allows Large Language Models (LLMs) to leverage external tools for solving complex user tasks, has emerged as a promising avenue for extending model capabilities. However, current approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model. In this paper, we propose ToolACE-R, a novel method that introduces adaptive self-refinement for tool invocations. Our approach features a model-aware iterative training procedure that progressively incorporates more training samples based on the model’s evolving capabilities. Additionally, it allows LLMs to iteratively refine their tool calls, optimizing performance without requiring external feedback. To further enhance computational efficiency, we integrate an adaptive mechanism when scaling the inference time, enabling the model to autonomously determine when to stop the refinement process. We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models, even without any refinement. Furthermore, its performance can be further improved efficiently through adaptive self-refinement. Our results demonstrate the effectiveness of the proposed method, which is compatible with base models of various sizes, offering a promising direction for more efficient tool learning.
ToolACE-R: Tool Learning with Adaptive Self-Refinement
Xingshan Zeng1, Weiwen Liu1, Xu Huang2, Zezhong Wang3, Lingzhi Wang4, Liangyou Li1, Yasheng Wang1, Lifeng Shang1, Xin Jiang1, Ruiming Tang1, Qun Liu1 1Huawei Noah’s Ark Lab 2University of Science and Technology of China 3The Chinese University of Hong Kong 4Harbin Institute of Technology, Shenzhen zeng.xingshan@huawei.com, liuweiwen8@huawei.com, xuhuangcs@mail.ustc.edu.cn
1 Introduction
Tool learning, which enables Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention. With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans Qu et al. (2024b). While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools Qin et al. (2023); Patil et al. (2023); Liu et al. (2024b, c).
Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models Liu et al. (2024b, c); Wang et al. (2024b); Abdelaziz et al. (2024). However, data synthesized by advanced models from different sources can lead to compatibility issues. Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations Ren et al. (2024); Kang et al. (2024). As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
Additionally, the potential of maximizing the model’s own capabilities remains under exploration in the field of tool learning. It has been shown that a model’s reasoning performance can be improved by scaling test-time computation Brown et al. (2024); Snell et al. (2024). One promising approach involves iteratively refining the model’s output multiple times and then selecting the majority answer Snell et al. (2024). However, user queries vary in complexity, ranging from simple to highly complex. Scaling test-time computation for overly simple and straightforward questions is inefficient, as models can often answer them correctly without additional refinement. Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of effort required for each question.
To address the challenges outlined above, we propose ToolACE-R, a method that enables adaptive self-refinement of tool invocations through a customized iterative training procedure. Specifically, we introduce a model-aware iterative training approach that facilitates LLMs in learning self-refinement for tool invocation in an iterative fashion, enhanced by model-based data selection. Additionally, by constructing refining data samples that preserve identical cases, LLMs are more likely to learn appropriate stopping criteria during the iterative process. Building on this training procedure, we further integrate adaptive self-refinement into iterative inference for inference time scaling, allowing models to autonomously determine when to halt the iteration, thereby improving scaling efficiency.
We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) Yan et al. (2024) and API-Bank Li et al. (2023), to assess the effectiveness and efficiency of ToolACE-R.The experimental results show that ToolACE-R achieves competitive performance compared to advanced API-based models like GPT-4o, and can be further enhanced efficiently through adaptive self-refinement. Additional analyses highlight the impact of the proposed modules and examine the generalizability of the method across different model backbones and sizes.
The key contributions of this work are summarized as follows:
  • We propose a model-aware iterative training procedure, enabling LLMs learning adaptive self-refinement for tool learning, which maximizes the model potential without any feedback.
  • We integrate adaptive self-refinement into iterative inference, enabling models to autonomously decide when to halt the refinement process and allowing for more efficient use of computational resources when scaling test-time compute.
  • We conduct extensive experiments on several representative tool calling benchmarks, revealing effectiveness and efficiency of our method.
2 Related Work
Tool Learning.
Integrating external tools can enhance LLMs’ capabilities, enabling more specialized, precise, and reliable problem-solving for users (Qin et al., 2023). Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches. Tuning-free methods employ prompting techniques, such as in-context tool-use examples, without additional training Mialon et al. (2023); Hsieh et al. (2023); Ruan et al. (2023). A notable example method is ReAct Yao et al. (2023), which can alternate between reasoning and action for complex tasks. Recent work Qu et al. (2024a) improves tool understanding by iteratively refining documentation. Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention Qin et al. (2023); Schick et al. (2023); Patil et al. (2023); Tang et al. (2023); Abdelaziz et al. (2024). Due to data scarcity, a large amount of them focus on data synthesis for improved generation procedures Liu et al. (2024b, c); Wang et al. (2024b), often overlooking the model’s inherent potential. While some others integrate tool feedback for refinement Du et al. (2024); Wang et al. (2024a), our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
Data Selection.
Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs Albalak et al. (2024). When the base model is sufficiently powerful, a small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of general data Zhou et al. (2023); Liu et al. (2023). While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection Du et al. (2023); Li et al. (2024). This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance Ren et al. (2024). Our method aligns with this perspective, proposing the use of pass@k metric to identify and select suitable training samples based on the model to be trained.
Self Refinement.
Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback Madaan et al. (2023); Weng et al. (2023) or external feedback Qu et al. (2024c); Xu et al. (2024). Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously Huang et al. (2024), which means that scaling test-time compute relying on iterative refinement still often depend on post-processing techniques, such as self-consistency Snell et al. (2024). In this work, we propose a simple yet effective method that enables models to adaptively self-refine their outputs, thereby improving scaling efficiency.
Figure 1: The overview of ToolACE-R, which includes two components, for training and inference, respectively.
3 Methodology
3.1 Problem Formulation
Tool learning aims to address the following problem: given a user query $q$ and a set of candidate tools $T=\{t_{1},t_{2},\dots,t_{n}\}$, the objective is to generate the correct tool invocations. This includes selecting the most appropriate tools and extracting the suitable parameter values. Specifically, the goal is to produce:
$A=[t_{1}(a_{1}),\dots,t_{m}(a_{m})]=f_{\theta}(\langle q,T\rangle)$ (1)
where $t_{j}$ and $a_{j}$ represent the $j$-th invoked tool and its corresponding argument, respectively, with $1\leq j\leq m$ and $m$ being the total number of tool invocations needed. The function $f_{\theta}(\cdot)$ denotes the generation process of the LLM with parameters $\theta$.
The training sample for tool learning is typically formalized as a triplet consisting of a user query, candidate tools, and the ground-truth tool invocation: $\{\langle q,T\rangle,A\}$, where $q$ and $T$ serve as input and $A$ should be output.
Our ToolACE-R is trained with a model-aware iterative training pipeline and can iteratively self-refine during inference in an adaptive manner. Figure 1 shows the overall procedure.
3.2 Model-Aware Iterative Training
Our model training starts with an instructed base model, ensuring that it has basic instruction following capability for our model-aware module. For training, we first gather a set of off-the-shelf training samples. Instead of using them directly, we apply a selection criterion to curate the data, which is then combined with generated self-refinement data. This curated and refined data is integrated into an iterative training process.
Data Selection Criterion.
Inspired by self-improvement techniques in existing LLM reasoning research Zelikman et al. (2022), where LLMs are trained using their own outputs, we introduce a pass@k criterion to select the most suitable training samples for each base model – those the model can answer (almost) correctly within at most $k$ tries. This approach may seem counterintuitive, as one might expect models to focus on learning from samples they struggle with. However, recent research Ren et al. (2024) has highlighted that models are more likely to perform self-alignment rather than learning new information during the fine-tuning stage. Based on this insight, we adopt this pass@k metric for data selection.
Specifically, for each sample $\{\langle q,T\rangle,A\}$, the base model to be trained first generate the invocation up to $k$ times. If any of the responses approaches the groundtruth $A$, we retain it as an official training sample:
$R(\langle q,T\rangle,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (2)
Here $R(\langle q,T\rangle,A)=1$ indicates the training sample will be retained. $\theta_{0}$ represents the base model before training and $f_{\theta_{0}}(\cdot)_{j}$ means the $j$-th generation with the base model.
A key aspect is how to define when a response "approaches" $A$. In our preliminary experiments, we found that if we restrict them to be exactly equal, the resulting training set still has an effect but limits the potential for performance improvement (see Table 4). However, by loosening this constraint, more suitable training samples are retained and the model’s potential can be more effectively harnessed. Therefore, we use the concept of "approaching" in this context. For any two invocations $A$ and $A^{\prime}$, we define $A\approx A^{\prime}$ if the used tools are exactly the same and each invocation only contains at most 1 different parameter values.
Self-Refine Data Construction.
To establish the model’s self-refinement capability, we additionally construct self-refine data based on the training samples. Specifically, we create a self-refine sample as a multi-turn interaction, where the first turn is the initial response from the model, and the second turn includes a refinement prompt from the user followed by the model’s refined answer: $\{\langle q,T\rangle,A_{1},r,A_{2}\}$. We apply the same selection criterion as described above, retaining the self-refinement training sample only when $A_{2}\approx A$:
$R(\langle q,T\rangle,A_{1},r,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle,A_{1},r)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (3)
Here, we replace $A_{2}$ with $A$ to create the refined training sample: $\{\langle q,T\rangle,A_{1},r,A\}$.
Notably, there may be samples where $A_{1}=A$, meaning the refined answer is identical to the previous answer. We retain these samples to enable the model to learn that "no further changes are needed when the answer is perfect." This is a crucial aspect of establishing the model’s adaptive, iterative self-refinement, and we will provide further details in Section 3.3.
Iterative Training.
Given the selected training samples (including self-refine samples), we apply supervised fine-tuning to the base model for one epoch. This constitutes one iteration. In subsequent iterations, we treat the trained model as the new base model, re-construct the self-refine samples, and re-select those that are suitable for fine-tuning the updated model. The model is then further trained using the newly selected samples.
Ideally, with each iteration, the model’s performance improves, leading to an increase in the number of selected samples. The iteration process continues until the number of selected samples no longer increases, indicating that the model has saturated on this set of training samples. The resulting model will then be our final ToolACE-R model.
3.3 Adaptive Self-Refine Iterative Inference
With model-aware iterative training, our ToolACE-R model can directly generate tool invocations based on a user query and candidate tools, and refine its answer when necessary.
We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration. The details are introduced in Algorithm 1. Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A_{0}$. In each subsequent iteration, ToolACE-R leverages the answer from the previous iteration as context and refines it. As mentioned earlier, ToolACE-R has learned that when the answer is sufficiently accurate, it can produce the same answer in subsequent refinement. Therefore, the iteration continues until the answers from two consecutive iterations are identical. We refer to this procedure as adaptive self-refine, where the model itself determines when to stop the iterations. To prevent infinite iteration, we set a maximum iteration limit $n$.
Algorithm 1 Adaptive Self-Refine
0:  Model $\theta$, query $q$, tools $T$, refine prompt $r$, time $n$
0:  Refined final answer $A$
1:  $A_{0}=f_{\theta}(\langle q,T\rangle)$  // First get the direct answer $A_{0}$
2:  for $i=1$ to $n$ do
3:     $A_{i}=f_{\theta}(\langle q,T\rangle,A_{i-1},r)$  // Get the refined answer
4:     if $A_{i}==A_{i-1}$ then
5:        return  $A_{i}$  // Adative stopping
6:     end if
7:  end for
8:  return  $A_{n}$  // $n$ is the maximum iteration time
Through adaptive self-refine, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
4 Experiments
Non-Live Live Overall
Models Simple Multiple Parallel
Parallel
Multiple
Simple Multiple Parallel
Parallel
Multiple
Non-live Live Overall
GPT-4o-2024-11-20 79.42 95.50 94.00 83.50 84.88 79.77 87.50 75.00 88.10 80.75 84.43
Gemini-1.5-Pro-002 78.33 93.50 92.5 90.00 82.95 77.40 87.50 83.33 88.58 78.68 83.63
GPT-4o-mini-2024-07-18 80.08 90.50 89.50 87.00 81.40 76.73 93.75 79.17 86.77 77.87 82.32
GPT-3.5-Turbo-0125 77.92 93.50 67.00 53.00 80.62 78.63 75.00 58.33 72.85 78.61 75.73
DeepSeek-V3 78.67 95.5 91.00 91.50 83.72 82.15 81.25 62.50 89.17 82.09 85.63
Llama3.1-70B-Inst 77.92 96.00 94.50 91.50 78.29 76.16 87.50 66.67 89.98 76.53 83.26
Llama3.1-8B-Inst 71.50 93.50 86.50 86.00 73.26 68.95 56.25 50.00 84.37 69.28 76.83
Qwen2.5-7B-Inst 70.50 94.00 90.50 82.50 74.81 71.89 62.50 62.50 84.38 72.17 78.28
ToolACE-8B (FC) 76.67 93.50 90.50 89.50 73.26 76.73 81.25 70.83 87.54 76.02 81.78
Hammer2.1-7B (FC) 78.08 95.00 93.50 88.00 76.74 77.4 81.25 70.83 88.65 77.20 82.92
xLAM-7B-r (FC) 74.25 95.50 81.00 73.50 72.09 74.93 50.00 62.50 81.06 73.87 77.47
ToolACE-R (FC) 79.92 96.50 93.50 93.00 81.78 81.01 87.50 83.33 90.73 81.27 86.00
  + Adaptive Self-Refine 81.17 96.00 94.00 93.00 82.95 81.58 87.50 83.33 91.04 81.94 86.49
Table 1: Accuracy comparison on BFCL (Last updated on 2025-01-18). The best results in each category are marked in bold. The second best results are underlined. FC indicates the models are fine-tuned for function calling.
4.1 Experimental Settings
Datasets and Models.
We leverage a subset of ToolACE Liu et al. (2024b) training data as our original full training dataset. The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English. For base model to be trained, we select LLaMA3.1-8B-Instruct AI@Meta (2024) for our main experiments. We also conduct experiments on other base models, including Qwen2.5-Instruct-series Yang et al. (2024) (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 Jiang et al. (2023), to validate the generalizability of ToolACE-R. We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-1.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 Liu et al. (2024a), Llama3.1-8B/70B-Instruct AI@Meta (2024) and Qwen2.5-7B-Instruct Yang et al. (2024). We also compare with fine-tuned tool calling models like ToolACE-8B Liu et al. (2024b), Hammer2.1-7B Lin et al. (2024) and xLAM-7b-r Zhang et al. (2024).
Benchmarks and Evaluation.
We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) Yan et al. (2024), ACEBench Chen et al. (2025), API-Bank Li et al. (2023) and ToolAlpaca Tang et al. (2023). To fully explore the model’s potential while minimizing external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries. For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively. Each subset includes four categories: Simple, Multiple, Parallel, and Parallel Multiple. Simple and Multiple cases both involve a single invoked tool, but in Multiple cases, there are multiple candidate tools. Parallel and Parallel Multiple cases require invoking multiple tools simultaneously. For ACEBench, we evaluate only the English normal category, excluding multi-turn cases. This includes the test samples in atom category and single-turn category. For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback. Performance is reported in terms of accuracy for all benchmarks.
Implementation Details.
Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA Hu et al. (2022). All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1. For data selection, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs. Each sample is generated up to 8 times (pass@8) to ensure potential. For the constructed refined data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn. During evaluation, we adopt greedy search to ensure stability and impose a maximum iteration limit of $n=5$.
4.2 Main Results
Results on BFCL.
Table 1 presents the comparison results on BFCL, including detailed results for each category and the overall average performance. The following key observations can be made:
  • The performance gap between API-based and open-source models is minimal. Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT-4o and Gemini-1.5-Pro.
  • Specialized fine-tuned models significantly benefit from domain-specific fine-tuning, allowing smaller models (7/8B) to compete with larger general models like Llama3.1-70B-Inst.
  • Our model, ToolACE-R which leverages our proposed iterative training procedure, achieves the best overall performance, surpassing both large API-based and open-source models.
  • The adaptive self-refinement mechanism further enhances the performance of ToolACE-R, with larger improvement observed in more complex categories (Live), highlighting the effectiveness of our inference method.
Models ACEBench APIBank ToolAlpaca
GPT-4o-2024-11-20 87.00 77.16 83.87
GPT-4o-mini-2024-07-18 81.63 73.10 83.87
Llama3.1-8B-Inst 50.63 60.41 79.03
ToolACE-R 83.88 75.89 85.48
  + Adaptive Sef-Refine 84.00 81.22 88.71
Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks.
Results on More Benchmarks.
To provide a more comprehensive evaluation, we continue to conduct experiments on three other representative benchmarks ACEBench, API-Bank, and ToolAlpaca, focusing on single-turn tool-calling scenarios. The results are summarized in Table 2.
On ACEBench, ToolACE-R demonstrates an impressive absolute improvement of over 30% compared to the base model, Llama3.1-8B-Inst, highlighting the effectiveness of our training pipeline. However, the benefit of Adaptive Self-Refine is marginal, likely due to the presence of particularly challenging cases where the model struggles to correctly refine its answers.
For both API-Bank and ToolAlpaca, ToolACE-R achieves substantial improvements. Moreover, when combined with Adaptive Self-Refine, the performance sees further gains, even surpassing GPT-4o. These results validate the potential of Adaptive Self-Refine to progressively refine tool invocation answers based solely on the model’s own capabilities, opening up new possibilities for scaling inference in tool learning scenarios.
Models Non-live Live Overall
Llama3.1-8B-Inst (Base) 84.37 69.28 76.83
ToolACE-R (Full) 91.04 81.94 86.49
  - Adaptive Self-Refine 90.73 81.27 86.00
    - Iterative Training 88.12 80.24 84.18
      - Data Selection 88.96 81.13 85.04
        - Refine Data 88.75 80.09 84.42
Table 3: Ablation Study of our proposed modules.
Ablation Study.
To investigate the impact of the proposed modules, we conduct an ablation study by progressively removing them. The corresponding evaluation results on the BFCL dataset are shown in Table 3. As observed, removing both adaptive self-refinement and iterative training leads to a noticeable performance degradation. However, further removing data selection does not result in a continued drop in performance. In fact, performance improves. This suggests that model-aware data selection should be combined with iterative training to fully leverage the entire training set, as the initial data selection (performed by the base model) typically only captures a small subset of the available data (mostly easy ones). Iterative selection is needed to capture more complex cases, as the trained model is improved. Finally, removing the constructed refine data also negatively impacts performance, validating that self-refine data acts as a form of data augmentation, also enhancing the model’s direct output.
4.3 Effects of Iterative
In this subsection, we examine the effects of our two iterative modules, i.e., model-aware iterative training and adaptive self-refine iterative inference.
Figure 2: Performance of each iteration during training. “Amount” indicates the amount of training samples in that iteration with our data selection module. Iteration 0 means the model before training, i.e., Llama3.1-8B-Inst.
Iterative Training.
Figure 2 illustrates the results of the iterative training process, showing both the accuracy performance and the amount of training data selected at each iteration. Iteration 0 corresponds to the base model, Llama3.1-8B-Inst. As depicted, the quantity of selected samples increases monotonically across iterations, suggesting that the model’s performance on the training set improves, as the selected samples are those the model is more likely to answer correctly (as indicated by pass@k). On the other hand, accuracy improves rapidly during the first two iterations, reaching its peak at iteration 3. Subsequently, accuracy plateaus, with no further improvement in later iterations. This suggests that excessive training could lead to overfitting on the training set.
Figure 3: Performance of different iterative self-refine methods. “Vanilla Self-Refine” always picks the final answer at each iteration. “Adaptive Self-Refine” is our proposed method, while “Self-Consistency” selects the majority answer during the iteration.
Iterative Inference.
To assess the effects of our adaptive self-refine method for iterative inference, we conducted additional experiments comparing it with two other inference-time scaling approaches. The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer. The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n. The results of these inference methods on BFCL, all based on the same model, are shown in Figure 3.
As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly. In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations. The "Self-Consistency at Itr n" method likely represents an upper bound for answer selection, with our method achieving competitive performance compared to this strong baseline. Furthermore, the performance of our method remains consistent across different iteration counts (3, 5, 10), likely because most cases terminate early, well before reaching the maximum iteration limit. This observation is supported by the average iteration time of 2.4, as indicated by our statistics on BFCL.
4.4 Effects of Data Selection
In this subsection, we investigate the effects of our model-aware data selection module. First, we explore the impact of the $k$ value when using pass@k for sample selection. Next, we examine how the use of data selected by different models affects model performance, highlighting the impact of model-aware selection.
Models Direct Refine Amount
Select with Pass@8
  First Iteration 80.24 80.61 120K
  Best Iteration 81.27 81.94 150K
Select with Pass@1
  First Iteration 78.90 78.53 94K
  Best Iteration 79.57 80.01 143K
Table 4: Accuracy performance on BFCL Live category when data selection with Pass@8 and Pass@1. “Direct” and “Refine” indicate the results with direct output and adaptive self-refine, respectively. “Amount” corresponds to the amount of selected training samples.
Data Selection with Pass@k.
In our data selection criterion, we retain training samples that the base model can answer (approximately) correctly within at most $k$ tries (pass@k). In this section, we examine the effects of varying $k$ values. Specifically, we report the evaluation results using pass@8 and pass@1 in Table 4. As shown, although the final number of selected samples is similar between the two methods (150K vs. 143K), the accuracy performance for pass@8 is considerably higher. The performance gap is already pronounced in the first iteration, suggesting that a more lenient selection criterion plays a crucial role in maximizing model potential and shaping the model’s distribution. Notably, this effect does not diminish over additional iterations. Based on these findings, we select a larger $k$ value for our experiments.
Model-aware Selection.
In our data selection process, we propose selecting training samples in a model-aware manner, ensuring that a sample is chosen only when the base model can answer it (almost) correctly. This is based on the assumption that samples aligning with the model’s distribution are more suitable for training. To further assess the impact of such a model-aware approach, we conduct an additional experiment. Instead of using data selected by the base model Llama3.1-8B-Inst, we train Llama3.1-8B-Inst with data selected by Qwen2.5-3B-Inst and Qwen2.5-7B-Inst, as well as with the intersection of all three models’ selected data – i.e., only the samples that all three models perform well on are included in the dataset. The accuracy results of the trained models, along with the corresponding training data amounts, are presented in Table 5. As shown, performance drops when using data selected by other models, even when the data volume is larger. The intersection dataset, which contains less data, achieves comparable results. This observation underscores the benefits of model-aware data selection.
Models Accuracy Amount
ToolACE-R 80.61 120K
Data selected with Qwen2.5-3B 79.50 125K
Data selected with Qwen2.5-7B 80.38 134K
Intersection of Three Datasets 80.27 108K
Table 5: Accuracy performance on BFCL Live category in the first interation when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
4.5 Further analysis
In this subsection, we examine the generalizability of our method by varying the base model and exploring the scaling effects across different model sizes.
Models Non-live Live Overall
Raw (Qwen2.5-7B-Inst) 84.38 72.17 78.28
  + Direct Data 88.29 78.39 83.34
  + ToolACE-R 89.23 81.20 85.21
Raw (Mistral-7B-Inst-v0.2) 55.25 55.07 55.16
  + Direct Data 84.06 70.17 77.12
  + ToolACE-R 87.98 80.01 84.00
Table 6: Accuracy performance of Different backbones on BFCL. “+ Direct Data” indicates the model directly trained with the full original set, while “+ ToolACE-R” is the result with our method, both iterative training and adaptive self-refine are applied.
Different Backbones.
To assess the generalizability of ToolACE-R, we conduct experiments with different base models, including Qwen2.5-7B-Inst and Mistral-7B-Inst-v0.2. The results, along with the original performance and performance after training with the full original training set, are presented in Table 6. As shown, fine-tuning with the full training set leads to a significant performance boost for both models, particularly for the relatively weaker Mistral-7B-Inst-v0.2, which sees an improvement of over $20\%$. Moreover, our method further enhances performance, using exactly the same original data, demonstrating its effectiveness.
Model Size Scaling.
Figure 4: Performance of models in Qwen2.5-Series. “Base” refers to the Qwen2.5-Instruct models, while “Direct” and “Adaptive Self-Refine” represent the results of models trained using our method and evaluated with direct output and Adaptive Self-Refine, respectively.
We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases. For this, we use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models. The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R. The direct output of ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B, and the Adaptive Self-Refine inference technique further enhances these results. This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
5 Conclusion
In this paper, we introduce ToolACE-R, a tool learning technique designed to enable large language models (LLMs) to effectively leverage external tools. ToolACE-R consists of two key components: model-aware iterative training and adaptive self-refine iterative inference. Model-aware iterative training maximizes the model’s potential by iteratively selecting the most appropriate samples for training. Adaptive self-refine further enhances performance by allowing the model to autonomously determine when to stop refining its own answers during iterative inference. Experimental results demonstrate that ToolACE-R consistently improves performance across various tool-calling benchmarks and is effective with different backbone models. Additionally, our analyses highlight the impact of the proposed modules on model performance.
Limitations
We summarize the limitations in the following two points:
First, due to resource constraints, we conduct experiments only on models with up to 8B parameters, using LORA. As shown in the scaling performance results in Figure 4, we have not yet observed any bottlenecks preventing larger models from benefiting from our ToolACE-R method.
Second, our focus is on single-turn tool-calling scenarios, excluding the multi-turn and multi-step cases for simplicity. However, we believe that single-turn cases are a fundamental aspect of tool learning. The significant improvements observed across several representative benchmarks demonstrate the effectiveness of our method. We plan to explore ToolACE-R in more complex scenarios in future work.
Appendix A Appendix
A.1 Detailed Introduction to Benchmarks
Below we introduce the used benchmarks with more details:
  • BFCL Yan et al. (2024): BFCL has progressed over time and is currently at version 3, which includes three main categories: non-live, live, and multi-turn. This paper focuses on the non-live and live categories, which primarily consist of Python-style tool call data. These are further subdivided into four types: Simple, Multiple, Parallel, and Parallel Multiple.
    • Single: A single evaluation is the simplest and most frequently encountered format, where the user provides a single function document, triggering exactly one tool call.
    • Multiple: The multiple category involves a user query that selects a single tool call from 2 to 4 available function documents. The model must choose the most appropriate tool to call based on the user’s context.
    • Parallel: A parallel test sample involves triggering multiple tool calls simultaneously in response to a single user query. The model must determine how many calls are needed.
    • Parallel Multiple: The parallel multiple cases combine the concepts of both parallel and multiple. Here, the model is provided with several function documents, and each corresponding tool call may be invoked zero or more times.
  • ACEBench Chen et al. (2025): ACEBench is a comprehensive tool-use benchmark that offers more detailed granularity. It is categorized into three main types: Normal, Special, and Agent. In this work, we focus on the Atom and Single-turn sub-categories within the Normal category. Atom cases involve a set of APIs that contain specific parameter types, such as booleans, enumerations, numbers, lists, and objects. The Single-turn category includes both single and parallel cases.
  • API-Bank Li et al. (2023): API-Bank is a dialogue-style tool call dataset, consisting of two settings: Call and Retrieve + Call. In this dataset, the model is tasked with invoking predefined local Python tools based on the user’s requirements in the dialogue. Accuracy is assessed by comparing whether the tool’s returned values match the ground truth. In this work, we focus on the Call setting in our experiments.
  • ToolAlpaca Tang et al. (2023): It introduces a multi-agent simulator that uses GPT-4 to simulate the return values of tools. The model can make modifications and re-call the tool based on the returned values (e.g., error messages). Ultimately, GPT-4 is used to evaluate the accuracy. In our experiments, we does not leverage the returned values for refinement for consistency.
A.2 Detailed Results in ACEBench
Table 7 displays the detailed results in each sub-category of ACEBench. As can be seen, in all sub-categories, ToolACE-R performs much better than its base model Llama3.1-8B-Inst. The improvement of Adaptive Self-Refine is not significant.
Atom Single-Turn Overall
Models Bool Enum Number List
Object
Short
Object
Deep
Single Parallel Overall
GPT-4o-2024-11-20 94.00 94.00 98.00 94.00 64.00 96.00 84.00 72.00 87.00
GPT-4o-mini-2024-07-18 88.00 94.00 98.00 90.00 44.00 92.00 77.00 70.00 81.63
Llama3.1-8B-Inst 50.00 74.00 48.00 60.00 30.00 44.00 55.00 44.00 50.63
ToolACE-R 88.00 96.00 94.00 94.00 50.00 92.00 83.00 74.00 83.88
  + Adaptive Self-Refine 88.00 96.00 94.00 94.00 50.00 94.00 84.00 72.00 84.00
Table 7: Detailed accuracy comparison on ACEBench.
A.3 Prompts
We show the detailed system prompt and refine prompt in Figure 5 and Figure 6, respectively. Note that the displayed prompts are those used during inference. For the training data, we diversify the system prompts to ensure better generalizability.
A.4 Data example
Figure 7 is an example of our refine data. The model will only learn the last turn, with previous turns as context.
You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections. If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke: {candidate tools} {other information}
Figure 5: The system prompt for tool calling. “{candidate tools}” is replaced with actual tool descriptions. “{other information}” is replaced with some other information in the case, e.g., date. If no other information, it is null.
Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect.
Figure 6: The refine prompt for tool calling.
[System] You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections. If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
    "name": "bookFlight",
    "description": "Book a flight for a specified destination",
    "parameters": {
        "type": "dict",
        "properties": {
            "origin": {
                "type": "string",
                "description": "The departure airport or city"
            "destination": {
                "type": "string",
                "description": "The destination airport or city"
            "departure_date": {
                "type": "string",
                "description": "The date of departure (in the format of YYYY-MM-DD)"
        "required": [
            "origin",
            "destination",
            "departure_date"
The current time is 2024-10-02 18:18:11. [User] I need to book a flight from New York to London for a business trip tomorrow. [Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]] [User] Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect. [Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]]
Figure 7: One example for refine data. In this example, the assistant does not change the tool invocation as it is correct already. Through this kind of data, the model is able to learn when to stop iteration during inference.